# Dynamic Intelligence Assessment: Benchmarking LLMs on the Road to AGI with a Focus on Model Confidence

### v1

@@ IV-A An Enthusiasm for Mathematics
One of the most striking observations is GPT-4o API’s eagerness to engage with mathematical tasks, even though it consistently fails to solve them correctly, as shown in Figure 6. This behaviour could stem from the fact that GPT-4o API and ChatGPT-4o share the same underlying architecture. In mathematics, GPT-4o consistently assumes that it has the capacity to solve the problems. This aligns with the observation made by Mirzadeh et al. [14], where large language models (LLMs) often rely on sophisticated pattern matching rather than true logical reasoning.
GPT-4o API’s inability to skip tasks reflects a broader issue in its self-assessment. Despite prompts encouraging it to skip problems it cannot solve, GPT-4o API rarely does so. This indicates a lack of meta-cognitive awareness in gauging its limitations, particularly when tools are required to solve a task. It continues to attempt math problems, even though it lacks the necessary tool-using capabilities to succeed.
Figure 6: Comparison of (a) GPT-4o results and (b) ChatGPT-4o results in Cybersecurity (100 templates, 500 questions generated) and Mathematics with (50 templates, 250 questions generated)
IV-B Cybersecurity vs. Mathematics: Different Approaches in Skipping
The contrast between the models’ behavior on mathematical tasks versus cybersecurity tasks provides additional insights. GPT-4o API, which eagerly attempts mathematical problems, is far more reserved in cybersecurity tasks, skipping questions almost 20 times more frequently (3 vs. 59 as shown in Figure 6). In cybersecurity tasks, both models tend to skip more frequently, suggesting that these tasks invoke a higher degree of self-reflection and caution. This behaviour indicates that both models are capable of critical evaluation, particularly when the task is complex enough to warrant skipping. Overall, ChatGPT-4o demonstrates an improved ability to assess when a task exceeds its capabilities, as evidenced by its higher tendency to skip difficult tasks.
@@ IV-C Limitations in Current Metrics
In the traditional setting, skipping is an inferior approach to hallucination or guessing. To underline this, we should examine ChatGPT-4o’s performance using a traditional metric like Pass@k. On the DIA-Bench, ChatGPT-4o achieved an impressive 73.3% Pass@5, meaning it was able to provide a correct answer to $73.3\%$ of the question templates when given five attempts. At the same time it was only able to ace all five instances for $38.67\%$ of the templates. The same goes for GPT-4o API, with Pass@5=21.3% and Conf@5=10.6%. The significant gap between the Pass@k score and the Confidence Index highlights the importance of reliability metrics.
Moreover, if we would rely on exact match for $k=1$, as is common with static datasets, at least a third of the tasks on DIA-Bench would have a minimum of 1 in 5 chance of being answered correctly through trial and error. This artificially boosts the success rate but does not accurately reflect a model’s reliability and confidence.
@@ IV-D A lack of Confidence
ChatGPT-4o recorded a high near-miss score, achieving 4 out of 5 on 14 templates. We anticipate that as $k$ increases, the near-miss score would rise, lowering the confidence index. During early experimentation with prompt templates, we observed that models often approached tasks with varying strategies. For example, when tasked with adding two smaller numbers, ChatGPT-4o might not immediately write and execute Python code—considering the task trivial—but instead produce an answer close to the correct one. For more complex tasks, it demonstrated this behaviour less.
ChatGPT-4o completely skipped 9 out of 150 templates, but inconsistently skipped at least one instance in 32 other templates, revealing gaps in its decision-making. This suggests that its self-assessment mechanism isn’t fully reliable, as it sometimes engages with tasks it later recognizes as too difficult. Despite this, ChatGPT-4o outperforms its API counterparts, which is expected since API models lack tools and code execution capabilities. Interestingly, both GPT-4o and ChatGPT-4o exhibit similar “taste” when attempting task categories, with GPT-4o taking on math problems enthusiastically despite lacking tools. In cybersecurity tasks, GPT-4o and ChatGPT-4o display a strong correlation in their skipping patterns—when GPT-4o skips a template, ChatGPT-4o often does the same for at least some instances of that template.
Other API models rarely utilized the skip option and attempted to answer every question. This pattern, evident in Table I, underscores that most models are primarily engaged in pattern matching and remain distant from achieving artificial general intelligence (AGI). The results reveal a substantial gap in problem-solving performance between models with tool-using capabilities and those without. LLMs without tools tend to hallucinate more often, failing to complete tasks accurately, which is reflected in their lower reliability scores.
@@ V Limitations and Threats to Validity
Certain task templates might not be as well represented as others, leading to an incomplete evaluation of the LLM’s overall performance. In our study, we focused on computer science and mathematics, omitting other disciplines to concentrate our resources. We aim to expand this, and the used data modalities.
We did not utilize o1-preview or other chat models due to their high time and cost requirements, which would have impacted the feasibility of this work.
@@ VI Conclusion
In this study, we introduce the Dynamic Intelligence Assessment (DIA) framework along with the DIA-Bench dataset to test the problem-solving reliability and confidence of large language models (LLMs) for mathematics and computer science questions. We also identified four key metrics for evaluating LLMs reliability and confidence. Our work addressed three key research questions:
@@
    Answer: By using hand-crafted dynamic question templates, hosting simpler tasks to very difficult ones across various formats (text, PDFs, compiled binaries, etc.). This, combined with the use of the proposed evaluation metrics, offers a more rigorous comparison of models’ abilities compared to static, non-multimodal, or overly easy benchmarks.
  • RQ3: In problem-solving, does the use of tools impact the confidence and reliability of different models?
    Answer: Yes, the use of tools significantly enhances both the confidence and reliability of models in problem-solving. Tool-using models like ChatGPT-4o not only perform better on complex tasks requiring precise computations, but they also demonstrate more consistent and reliable results across various task types. Interestingly, there is a strong correlation between the GPT models in their tendency to attempt to solve a task, where non-tool users are more prone to errors and hallucinations.
Our findings demonstrate that while current LLMs have made significant progress, there remain challenges in achieving consistent and reliable problem-solving, particularly for models without tool-using capabilities. DIA-Bench provides a valuable resource for further improving LLM evaluation and development towards more adaptable and reliable AI systems.

### v3

@@ IV-A An Enthusiasm for Mathematics
One notable observation is the persistent eagerness of most non-tool-using API models to engage with mathematical tasks, despite their consistent inability to solve them correctly, as illustrated in Figure 6. This behavior suggests a lack of meta-cognitive awareness, particularly in acknowledging their limitations with tasks requiring tool usage. Despite prompts encouraging these models to refrain from attempting problems they cannot solve, they rarely do so, resulting in unsuccessful attempts at mathematical problems.
A detailed examination of OpenAI’s models reveals that GPT-4o, GPT-4o-mini, and ChatGPT-4o exhibit similar levels of enthusiasm in attempting such tasks, possibly due to their shared underlying architecture. While GPT-4o and GPT-4o-mini consistently fail at mathematical problems due to the lack of code execution ability, ChatGPT-4o demonstrates proficiency in solving these questions.
This observation aligns with findings by researchers at Apple [14], noting that LLMs mostly rely on sophisticated pattern matching rather than genuine logical reasoning. As shown in Figure 6 and Table I, most models are incapable of skipping tasks, even when these tasks are infeasible to solve without tools. However, the dataset in [14] uses “relatively simple grade-school math questions, requiring only basic arithmetic operations”, and doesn’t allow non-tool using models to skip.
However, using the DIA framework, OpenAI’s o1-mini model presents a completely different picture. When given the option to skip questions, o1-mini conducts an assessment and refrains from attempting to answer tasks it deems beyond its reach, demonstrating significant meta-cognitive awareness in identifying which tasks to avoid. It is important to note that in 52 instances, o1-mini attempted to skip a question but failed to answer I-DO-NOT-KNOW in the specified format, resulting in the response being marked as incorrect (this issue is discussed further in Section V). These instances indicate that o1-mini’s capabilities may be more robust than reported. However, it achieves a high NMS, score ($4/5$ on 14 templates), indicating that its reliability is far from perfect.
IV-B Cybersecurity vs. Mathematics
The difference in model behavior between mathematical and cybersecurity tasks offers interesting insights. API models without tool-using capabilities, which readily attempt mathematical problems, show more restraint with cybersecurity tasks, as illustrated in Figure 6. Notably, GPT-4o skips questions nearly 20 times more often in the cybersecurity domain (3 vs. 67). Similarly, models like Mistral-Nemo or GPT-4o-mini, which never skipped mathematical tasks, skipped cybersecurity tasks 15 and 29 times, respectively. This suggests that these more complex tasks invoke a higher degree of self-reflection and caution. This behavior indicates that some models are capable of critical evaluation, particularly when the task is complex enough to warrant skipping. For certain cybersecurity prompts, some models objected to answering, which category is subject to False Refusal Rate, however measuring this was outside our focus.
@@ IV-C Limitations in Current Metrics
In the traditional setting, skipping is an inferior approach to hallucination or guessing. To underline this, we should examine ChatGPT-4o’s performance using a traditional metric like Pass@k. On the DIA-Bench, ChatGPT-4o achieved an impressive 73.3% Pass@5, meaning it was able to provide a correct answer to $73.3\%$ of the question templates when given five attempts. At the same time, it was only able to ace all five instances of $38.67\%$ of the templates. This artificially boosts the success rate, but does not accurately reflect a model’s reliability and confidence. A similar gap exists between most other well performing models, as can be seen in Table II.
Model Pass@5 (%) Conf@5 (%)
ChatGPT-4o 73.3 38.67
GPT-4o API 34.0 17.33
o1-mini 48.0 21.34
TABLE II: Performance metrics for ChatGPT-4o, GPT-4o API, and o1-mini.
The significant gap between the Pass@k score and the Confidence Index highlights the importance of reliability metrics. At the same time, if someone only focuses on the number of correct solutions they would be led to believe that ChatGPT-4o is a superior model to o1-mini. However, this is the result of ChatGPT capitalizing on the mathematics tasks where it was able to use tools. Regardless, o1-mini beats its tool equipped colleague in achieving the best Reliability Score.
@@ IV-D A Lack of Confidence
ChatGPT-4o recorded a high Neaer Miss Score, achieving 4 out of 5 on 14 templates. We anticipate that as $k$ increases, NMS would rise, lowering the confidence index. During early experimentation with prompt templates, we observed that models often approached tasks with varying strategies. For example, when tasked with adding two smaller numbers, ChatGPT-4o might not immediately write and execute Python code—considering the task trivial—but instead produce an answer close to the correct one. For more complex tasks, it demonstrated this behavior less. The highest Neaer Miss Score belongs to GPT-4o, achieving $4/5$ in 29 templates. This again highlights the risks of traditional metrics like Pass@k.
ChatGPT-4o completely skipped only 9 out of 150 templates, but inconsistently skipped at least one instance in 32 other templates, revealing gaps in its decision-making. This suggests that its self-assessment mechanism isn’t fully reliable, as it engages with tasks that later recognizes as too difficult.
In consistent skipping, o1-mini takes the lead with $23$ question templates completely skipped, GPT-4o managed $7$, while GPT-4o-mini was only able to stay consistent for $4$ templates. Despite this, ChatGPT-4o outperforms its API counterparts, which is expected since API models lack tools and code execution capabilities. Interestingly, both GPT-4o and ChatGPT-4o exhibit similar “taste” when attempting task categories, with GPT-4o taking on math problems enthusiastically despite lacking tools. In cybersecurity tasks, GPT-4o and ChatGPT-4o display a strong correlation in their skipping patterns—when GPT-4o skips a template, ChatGPT-4o often does the same for at least some instances of that template.
Other API models rarely utilized the skip option and attempted to answer most questions. This pattern, evident in Table I, underscores that most models are primarily engaged in pattern matching and remain distant from achieving artificial general intelligence (AGI). The results reveal a substantial gap in problem-solving performance between models with tool-using capabilities and those without. LLMs without tools tend to hallucinate more often, failing to complete tasks accurately, which is reflected in their lower reliability scores.
Most importantly, we observe that the mere availability of tools does not enhance a model’s overall judgment. ChatGPT-4o and GPT-4o exhibit similar performance and a comparable tendency in attempting tasks. However, o1-mini, despite lacking tools, surpasses both in reliability and self-reflection.
@@ V Limitations and Threats to Validity
Certain task templates might not be as well represented as others, leading to an incomplete evaluation of the LLM’s overall performance. In our study, we focused on computer science and mathematics, omitting other disciplines to economize our resources. We aim to expand this, and the used data modalities.
We did not utilize o1-preview or other chat models due to their time and cost requirements, which would have impacted the feasibility of this work.
[…]
@@ VI Conclusion
In this study, we introduced the Dynamic Intelligence Assessment (DIA) framework along with the DIA-Bench dataset to test the problem-solving reliability and confidence of Large language Models (LLMs) for mathematics and computer science questions. We also introduced four new key metrics for evaluating LLMs’ reliability and confidence. Our work addressed three key research questions:
@@
    Answer: By using dynamic question templates, hosting simpler tasks to very difficult ones across various formats (text, PDFs, compiled binaries, etc.). This, combined with the use of the proposed evaluation metrics, offers a more rigorous comparison of models’ abilities compared to static, non-multimodal, and overly easy benchmarks, with metrics like Pass@k or Accuracy.
  • RQ3: How does the use of tools impact the confidence and reliability of various models in problem-solving?
    Answer: While the use of tools enhances a model’s problem-solving ability, it doesn’t necessarily improve its judgment or self-assessment ability. GPT-4o and ChatGPT-4o share the same underlying architecture, and ChatGPT-4o performs better primarily because it has access to tools, not because of better chain-of-thought type reasoning. GPT-4o often attempts tasks it has no chance to solve, showing a lack of judgment. In contrast, o1-mini, a non-tool-using API model, demonstrates superior reasoning by effectively deciding when not to attempt unsolvable tasks, leading to a higher Reliability Score than the tool-using ChatGPT-4o. Therefore, while tools improve problem-solving performance, the model’s design and architecture are the critical factors determining confidence and reliability.
Our findings demonstrate that while current LLMs have made significant progress, there remain challenges to achieve reliable problem-solving capabilities. OpenAI’s o1 model family is to this date the first among the tested models, that exhibit capabilities beyond just simple pattern matching, and demonstrated excellent critical thinking and self-reflection.
Acknowledgement
This research is supported and funded by the Technology Innovation Institute (TII), Abu Dhabi. Additional support is provided by the TKP2021-NVA Funding Scheme under Project TKP2021-NVA-29; ELTE-OTP Cyberlab—a collaboration between Eötvös Loránd University (ELTE) and OTP Bank Plc; EPSRC grant EP/T026995/1 titled “EnnCore: End-to-End Conceptual Guarding of Neural Architectures” under the Security for All in an AI-enabled Society program; the Research Council of Norway Project No. 312122 “Raksha: 5G Security for Critical Communications”; funding from Horizon Europe under Grant Agreement No. 101120853; and funding under Grant Agreement No. 101145874, supported by the European Cybersecurity Competence Centre.